OpenStudy (anonymous):

write a slope-intercept equation of the line whos graph is parallel y=5x -3 and has the y intercept (0,7)

OpenStudy (anonymous):

If two lines are parallel, then they have the same slope.

OpenStudy (anonymous):

slope-intercept form is as follows: y = mx + b where m is the slope of the line and b is the y-intercept

OpenStudy (anonymous):

you could also use point slope form: y - y1 = m(x - x1) where y1 = 7 and x1 = 0